40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 29 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  15212   Fri Feb 14 00:53:50 2020 gautamUpdateALSFast ALS - more setup

In the process of setting up some cabling at 1Y2, I must've bumped a cable to the c1lsc expansion chassis. Anyways, the c1lsc models crashed. I ran the reboot script around 530pm PDT. Usual locking behavior was recovered after this. The work at 1Y2 was:

  • Ran a cable from X Beat power splitter ("LO" leg of the analog delay line) to variable delay line. 
  • Ran cable from delay line to demodulator's LO input.
  • Set up the SR785 for some CM board TF measurements.

The IN2 to CM board was already connected to I single ended output of the ALS X demodulator. The ~100 Hz UGF digital locking using the CM_SLOW path is straightforward but I didn't have any success with the AO path tonight. I wonder how high BW this lock can be made without injecting a ton of noise into the IMC loop, given that the EX uPDH only has ~ 10 kHz UGF.

Attachment #1 shows the spectra of the ALS signal 

  • The two "CM Slow" traces are the digitized "SLOW" output of the common mode board, whose IN2 is connected to the demodulated I output of the analog delay line.
  • The delay in the LO line of the analog delay line is adjusted to zero the DC value of this signal to best effort.
  • These spectra are measured with the arm cavities POX/POY locked, and the EX laser locked to the arm cavity using the end PDH box.
  • I simultaneously monitor the output of the digital phase tracker servo, and scale the CM Slow signal such that the spectra line up. The scaling factor required was to multiply the CM_SLOW signal x10 (CM board IN2 gain was set to +6dB, to account for the x2 gain in going from single ended to differential inside the ALS demodulator box).
  • One puzzline feature is why switching on the ADC whitening makes the ALS spectrum noisier (even though it clearly changes the digitization noise floor). There is a peak that appears at ~ 8 kHz with the whitening on, and it may also be downconverted noise from some peak at higher frequencies I guess (if the AA isn't sharp enough). 

Attachment #2 is an OLTF measurement.

  • In the blue trace, the arm length is controlled by using the CM Slow signal as an error signal, applying feedback to IMC length via MC2.
  • In the red trace, I turned the digital MC2 violin notches off, and added upped the IMC IN2 gain to -12 dB (AO gain slider = 0dB).
  • This was as high as I could go before the PC drive RMS began to go crazy.
  • But still, there isn't any significant phase advance.
  • It is possible I need to tack on a low-pass filter to prevent noise injection at higher frequencies...
Attachment 1: CMSlow_ALSnoise.pdf
CMSlow_ALSnoise.pdf
Attachment 2: OLTFmeas.pdf
OLTFmeas.pdf
  15211   Thu Feb 13 21:30:55 2020 shrutiUpdateALSALS OOL noise with arms locked

[Meenakshi, Gautam, Shruti]

Summary:

- We initially aligned the arm cavities to get the green lasers locked to them. For the X arm cavity, we tweaked the ITMX and ETMX pitch and yaw and toggled the X green shutter until we saw something like a TEM00 mode on the monitor and a reasonable transmitted power.

- With the LSC servo enabled, the IR light also became resonant with the cavities.

- Then we measured the noise in different configurations. Attachment 1 shows the the ALS OOL (in the IR beat signal) noise with the arms locked inidividually via PDH.


The script for plotting the ALS beat frequency noise is:

users/Templates/ALS/ALS_outOfLoop_Ref.xml
Attachment 1: 20200213_ALS.pdf
20200213_ALS.pdf
  15210   Thu Feb 13 02:07:26 2020 gautamUpdateLSCAO path transfer function measurement

Summary:

I measured the transfer function of the AO path, and think that there are some features indicative of a problem somewhere in the IMC locking loop.

Details:

Regardless of the locking scheme used, high bandwidth control of the laser frequency relies on the fact that the laser frequency is slaved to the IMC cavity length with nearly zero error below ~50 kHz (assuming the IMC loop has a UGF > 100 kHz). In my single arm experiments, I didn't know what to make of the ripples that became apparent in the measured OLTF as the AO gain was ramped up.

Tonight, I measured the TF of the "AO path", which modifies the error point of the IMC, thereby changing the laser frequency. 

  • An SR785 was used to make the measurement.
  • The signal was injected at the "EXC B" input on the CM board.
  • The CM_SLOW path was disabled, AO gain = 0dB, IMC IN2 gain = 0dB.
  • Between "EXC B" and the IMC error point (which I measured at TP1A on the IMC board), we expect that there are 2 poles at ~ 6 Hz, and one pole at ~ 11 Hz.

Attachment #1 shows the result of the measurement. 

  • This measurement should be the "Closed Loop Gain" [= 1/(1+L) where L is the open loop gain] of the IMC locking loop. For comparison, I've overlaid the inferred CLG from a measurement of the IMC OLG I made in Jun 2019. The magnitude lines up quite well, but the phase does not 🤔 
  • Above 10 kHz, the measurement is as I expect it to be.
  • However, between 1 kHz and 10 kHz, I see some periodic features every 1 kHz, which I don't understand. In the IMC OLTF, these would be sharp dips in the OLTF gain.
  • I was careful not to overdrive the servo, so I believe these features are not a measurement artefact.
  • Combing through past elogs, I couldn't really find any measurements of the IMC OLTF in the 1 kHz - 10 kHz band.
  • I decided to measure the spectrum of the IMC error point (with no excitation input), to see if that offered any additional insight. Attachment #2 shows the result - again, periodic features at ~ 1 kHz intervals.

I didn't use POX / POY as a sensor to confirm that this is real frequency noise, I will do so tomorrow. But it may be that realizing a stable crossover is difficult with so many features in the AO path.

Previous thread with a somewhat detailed characterization of the IMC loop electronics.

Attachment 1: AOpathTF.pdf
AOpathTF.pdf
Attachment 2: IMCinLoop.pdf
IMCinLoop.pdf
  15209   Thu Feb 13 01:47:39 2020 gautamUpdateALSFast ALS - delay line prep

A few years ago, Koji and I setup a delay line phase shifter, which can be used to impart a (switchable) delay to a signal path. Since we talked about reviving the fast (= high bandwidth) ALS control scheme at the meeting, I reminded myself of the infrastructure available.

  • Schematic
  • Comprehensive note on theory of operation / performance.
  • Past elog threads - #11603 and #11604.
  • Attachment #1 - my modification to the ALS screen to add a slider that controls the channel C1:LSC-BO_1_0_SET. The label is a bit misleading for now - elog11604 tells you the conversion between this slider value and the actual delay in nanoseconds, but I couldn't get a soft channel set up that correctly FLNKed to this record. In the process of trying to do so, I edited the C1_ISC-AUX_ALS.db file, and also restarted the modbus and latch processes on c1iscaux a few times.
  • Attachment #2 - frequency dependent loss for some representative delays. At ~200 MHz, I find the measured loss to be > 8dB, which is ~2dB more than what the D. Sigg note tells me to expect. This is rather a lot of loss, but I guess it's okay. Measurement cable loss was calibrated out with the AG4395A.
  • Attachment #3 - confirmation of constantness of delay as a function of frequency, for some representative delays. The "undelayed" setting corresponds to a fixed delay of ~4 nsec, which is consistent with what the D. Sigg note tells me to expect. Once again, I calibrated out the delay of the measurement setup using the AG4395A.

For a beat note in the regime 10-100 MHz, we should have plenty of range in this module to add a delay such that we zero one quadrature of the ALS DFD output (for a linear error signal). 

I then proceeded to connect the single-ended front panel BNC corresponding to the ALS_X_I DFD channel to the IN2 input of the CM board (this would be what we use for high bandwidth ALS feedback). The conventional ALS system uses the differential output from a rear-panel D-sub, so in principle, both systems could run in parallel. I confirmed that I could see a signal when the IN2 path on the CM board was engaged (monitored using ndscope at the CM_Slow output), and that this signal stabilized when the green laser was locked to the X-arm length, which itself was slaved to the PSL frequency using the usual POX locking scheme. I have not yet routed the LO leg of the ALS_X beat through the delay line phase shifter - see next elog for details.

Update about the ALS MEDM screen slider: the trick was to change the OMSL field of the C1:LSC-BO_1_0 channel to "closed_loop" instead of "supervisory". Once this is done, the DOL value of the same channel can be set to the soft channel C1:ALS-DelayCalc, which sets the 16 bit binary string that controls the delay. Because arbitrary delays are not possible, I think it's more natural for the user to interact with this 16-bit binary string rather than the actual delay itself. So the MEDM screen has been slightly modified from what is shown in Attachment #1.

Attachment 1: delaySlider.png
delaySlider.png
Attachment 2: delayLineLosses.pdf
delayLineLosses.pdf
Attachment 3: delayLineCal.pdf
delayLineCal.pdf
  15208   Wed Feb 12 12:13:37 2020 gautamUpdateLSCAO path attempts

Summary:

  1. The PRFPMI can be controlled by a mix of ALS and RF signals and circualting arm powers > 100 can be maintained for several tens of minutes at a stretch.
  2. The complete RF handoff still cannot be realized - I need to study the AO path crossover more carefully to understand what exactly is wrong and what needs to be done to rectify the problem.

Measurements:

Over the last couple of days, I've been trying to see if I can measure the phase advance due to the AO path - however, I've been unable to do so for any combination of CM board IN1 gain and MC Servo board IN2 gain I've tried. Yesterday, I tried to understand the loop shapes I was measuring a little more, and already, I think I can't explain some features.

Attachment #1 shows the TF measured (using SR785, and the EXC_A bank of the CM board) when the CM Slow path has been engaged.

  • All CARM control in this state is digital.
  • For the CM Slow path, the digital filter includes a pole at 700 Hz, pole at 5 kHz and zero at 120 Hz (the latter two for coupled cavity pole compensation).
  • In this conditions, the arm powers are somewhat stable at ~150, but still there are fluctuations of the order of 50%.
  • The "buzzing" as the arms rapidly go in and out of resonance is no longer present though.
  • The UGF of the hybrid REFL11+ALS loop is ~200 Hz, with ~45 deg of phase margin.
  • Turning off the MC2 violin filters gives some phase back. But I don't really understand the flattening of the TF gain between ~250-500 Hz.

Attachment #2 shows error signal spectra for the in-loop PRFPMI DoFs, for a few different conditions.

  • Engaging the REFL11 digital path smooths out the excess noise in the ~30-50 Hz band, which is consistent with the fact that the arm powers stabilize somewhat.
  • However, there is some gain peaking around ~400 Hz.
  • This is in turn imprinted on the vertex DoFs, making the whole system's stability marginal.

I believe that a stable crossover is hopeless under these conditions.

Next steps:

  1. Account for the measured OLTF, understand where the flattening in the few hundred Hz region is coming from.
  2. Repeat the high BW POY experiments, but with the simulated coupled cavity pole - maybe this will be a closer simulation to the PRFPMI transition.
Attachment 1: CARM_OLTF.pdf
CARM_OLTF.pdf
Attachment 2: PRFPMI_errSigs.pdf
PRFPMI_errSigs.pdf
  15207   Tue Feb 11 19:11:35 2020 shrutiUpdateComputer Scripts / ProgramsMATLAB on donatella

Tried to open MATLAB on Donatella and found the error:


MATLAB is selecting SOFTWARE OPENGL rendering.

 

License checkout failed.
License Manager Error -9
This error may occur when: 
-The hostid of this computer does not match the hostid in the license file. 
-A Designated Computer installation is in use by another user. 
If no other user is currently running MATLAB, you may need to activate.

Troubleshoot this issue by visiting: 
http://www.mathworks.com/support/lme/R2015b/9

Diagnostic Information:
Feature: MATLAB 
License path: /home/controls/.matlab/R2015b_licenses/license_donatella_865865_R2015b.lic:/cvs/cds/caltech/apps/lin
ux64/matlab15b/licenses/license.dat:/cvs/cds/caltech/apps/linux64/matlab15b/licenses/license_chiara_
865865_R2015b.lic:/cvs/cds/caltech/apps/linux64/matlab15b/licenses/license_pianosa_865865_R2015b.lic 
Licensing error: -9,57.


So I used  my caltech credentials to get an activation key for the computer. I could not find the option for a campus license so I used the individual single machine license.

Now it can be run by going to the location:

/cvs/cds/caltech/apps/matlab17b/bin

and running

./matlab

On opening MATLAB, there were a whole bunch of other errors including a low-level graphics error when we tried to plot something.

  15206   Tue Feb 11 16:39:00 2020 shrutiUpdateALSAM/PM

The results of the AM/PM measurements:

  • Attachment 1: Traces of 9 AM TFs overlaid on top of each other, calibrated by measuring the voltage at the ‘GREEN_REFL’ output where the TF was measured (described in elog 40m:15197). This was almost exactly 2 V.
  • Attachment 2: Traces of 9 PM TFs also overlaid measured using DLFD (as described in elog 40m:15180). Calibrated using the measured ~600 mV pk-pk voltage. The phase plots were unwrapped (shifted by 180 deg if needed) so that each started from roughly 0 deg.

Both the AM and PM TFs were scaled to make them have the same average value. Manually adjusting the delay line offset for each measurement using the oscilloscope was probably not accurate enough and therefore resulted in different scaling which this should somewhat compensate.

Attachment 3:

  • The orange and green lines are the averages of the PM and AM values of Attachments 1 and 2 respectively.
  • The solid red line is at 230 kHz, which was the previously chosen value for PDH locking. The peak seems to have shifted to the left from previous measurements (elog 40m:12077).
  • A horizontal black dashed line is drawn to show where the ratio is 10^5.
  • The red regions correspond to frequencies where PM/AM > 10^5 [only shown for frequencies greater than 200kHz], these are roughly (in kHz):
    • 211.4-213.9
    • 221.4-230.7 (peak at 225.642)
    • 240.8-257.9
    • ~748.3
    • 753.3-799.8, two largest peaks at 763.673 and 770.237
    • 809.6-829.3, peak at 819.472
    • 839.2-842.4
    • 881.8-891.7

Updated Calibration

Attachment 2 and 3 were miscalibrated due to an error in my understanding of the delay line, but the net result of the change in factors is qualitatively almost the same and the position of the major peaks remain predominantly unchanged.

The new plot is in Attachment 5.

The new calibration factor used: 5 MHz/V at the output of the mixer to obtain the frequency modulation and then division by the mod. freq. to obtain PM.

5 MHz/V because changing the PZT voltage by 0.01 V=> change in beat frequency by 0.1 MHz, which was seen as a 20 mV change in the delay line mixer output.

Again, the calibration is not very precise and I will probably repeat this experiment at some point more precisely.

Attachment 1: AM.pdf
AM.pdf
Attachment 2: PM.pdf
PM.pdf
Attachment 3: Ratio_all.pdf
Ratio_all.pdf
Attachment 4: Ratios_FM_PM.pdf
Ratios_FM_PM.pdf
Attachment 5: Ratio_all_new.pdf
Ratio_all_new.pdf
  15205   Mon Feb 10 15:55:46 2020 JordanUpdatePSLPMCTRANSPD

[Gautam, Jordan]

Gautam showed me how the PMCTRANSPD signal was reading zero, and he suspected it might have to do with the acromag wiring. Disconnected the acromag box underneath the PSL table and checked the ADC wiring. Side note: When benchtesting the c1psl acromag chassis there was excess noise in the AI channels, and grounding the minus pin of the ADC channel eliminates the noise.

So I grounded the (-) pins on the ADC1 (192.168.113.122), which PMCTRANSPD is connected to and that seemed to fix the problem. As of right now PMCTRANSPD is reading ~.75 V.

See attached pictures

gautam: While this fix seems to have worked, I wonder why this became necessary only in the last month. Note that the problem was a noisy readback on the PMC transmission PD, which also made the FSS_RMTEMP channel noisy, leading me to suspect some kind of ground loop issue.

Attachment 1: ADC1.jpg
ADC1.jpg
  15204   Mon Feb 10 15:54:47 2020 JordanUpdatePSLCompleted Acromag Wiring

All spare channels on the PSL acromag chassis are connected with ~12in of spare wiring for future use.

  15203   Mon Feb 10 15:04:42 2020 JordanUpdateGeneralHDMI Routing for new tv

Ran HDMI to the new tv mounted on the north wall of control room.

  15202   Mon Feb 10 10:07:20 2020 gautamUpdatePSLPMC re-locked

I found the PMC unlocked this morning. It was re-locked using the usual procedure. I feel like this has been happening more frequently in the last month than before. In the past, the cause seems to have been the PZT voltage drifting too close to one of the rails - however, in this case, it looks like an IMC unlock event is what triggered the PMC lockloss (admittedly the PZT voltage was somewhat close to the rail). It would be good if someone can re-connect the PMC Transmission photodiode, it was a useful diagnostic channel we had working fine before the ringdowns started.

I also tweaked the input pointing into the PMC and ran the WFS DC offset relief script.

Attachment 1: PMCunlock.png
PMCunlock.png
  15201   Mon Feb 10 09:40:54 2020 Larry WallaceSummaryGeneralSolidWorks Computer Upgrade and Printer repair

On February 5, 2020, the Dell engineering workstation located in the 40M lab, was replaced with a newer Engineering workstation, per a request from Koji . The new workstation should perform a good deal better over the older unit. It has more cores, more memory and a better video card. Since this unit is being used by the 40M group, the Comsol s/w pkg. was also installed on the unit.

During the computer swap, Koji had a problem with a print job and it was discovered the bottom tray of the HP5550 printer was broken. The broken tray was replaced from another unit that was being disposed of.

  15200   Fri Feb 7 19:39:10 2020 KojiUpdateLSCMore high BW POY experiments

This measurement tells you how the gain balance between the SLOW_CM and AO paths should be. Basically, what you need is to adjust the overall gain before the branch of the paths.

Except for the presence of the additional pole-zero in the optical gain because of the power recycling.

You have compensated this with a filter (z=120Hz, p=5kHz) for the CM path. However, AO path still don't know about it. Does this change the behavior of the cross over?

If the servo is not unconditionally stable when the AO gain is set low, can we just turn on the AO path at the nominal gain? This causes some glitch but if the servo is stable, you have a chance to recover the CARM control before everything explodes, maybe?

  15199   Fri Feb 7 15:00:16 2020 gautamUpdateLSCMore high BW POY experiments

To study the evilution of the AO path TFs a bit more, I've hooked up POY11_Q Mon to IN1 of the CM board. I will revert the usual setup later in the evening.

Update 1730: I've returned the cabling at 1Y2 to the nominal config, and also reverted all EPICS settings that I modified for this test. Y-arm POY locking works. Attachment #1 shows the summary of the results of this test - note that the AO gain was kept fixed at +5dB throughout the test. I have arbitrarily trimmed the length of the frequency vector for some of these traces so that the noisy measurement doesn't impede visual interpretation of the plots so much. At first glance, the performance is as advertised. I basically followed the settings I had here to get started, and then ramped up various gains to check if the measured OLTF evolved in the way that I expected it to. The phase lead due to the AO path is clearly visible.

Some important differences between this test and the REFL11 blending is (i) in the latter case, there will also be a parallel loop, CARM_A, which is effecting some control, and (ii) the optical gain of CARM-->REFL11_I is much higher than L_Y-->POY. So the initial gain settings will have to be different. But I hope to get some insight into what the correct settings should be from this test. I think IMC servo IN2 gain and AO gain slider on the CM board are degenerate in the effect they have, modulo subtle effects like saturation.

One possibility is that the gain allocation I used yesterday was wrong for the dynamic range of the CARM error signal. In some initial trials today, when I set the CM board IN1 gain to -32dB (as in the case of attempting the CARM RF handoff) and compensated for the reduced POY PDH fringe amplitude by increasing the digital gain for the CM_Slow path, I found that there was no phase advance visible even when I ramped up the IMC IN2 gain to +10dB. So, for the CARM handoff too, I might have to start with a higher CM board IN_1 gain, compensate by reducing the CM_Slow digital gain even more, and then try upping the IMC IN2 gain.

P.S. When the excitation input to the CM board was enabled in order to make TF measurements, I saw significant increase in the RMS of the error signal. Probably some kind of ground loop issue.

Attachment 1: AO_TFs.pdf
AO_TFs.pdf
  15198   Fri Feb 7 12:58:25 2020 YehonathanUpdateGeneralMetal PMC parts

I took the metal PMC box and examined its content and find the following items:

Name Quantity Picture (Attachment #)
Metal PMC body (PMC1) 1 1-3
Metal PMC body with two mounted 41 deg mirrors (PMC2) 1 4-6
"Baked PZT Caps" 3 7
PZT Caps 2 8
Flat mirror mounts 2 9
Bar clamps 4 10
Clamp studs 8 10
PZTs 4 11
ORings INF 12
Ball bearings INF 13
6.8 deg AoI curved mirrors (r=-1000mm) 6 14
41 deg AoI flat mirrors, R=99% @ 1064nm (1 Damaged) 11 15

There seem to be enough parts to build 2 PMCs + spares.

I find several problems in the metal PMCs:

PMC1 has a broken screw in one of its flat mirror mounts (Attachment 16). We need to get it out in the machine shop.

PMC2 one of the flat mirrors has a scratch on the AR coating and its ORing is failing (Attachment 17). Mirror and ORing need to be replaced.

I measure the physical dimensions of the PMC with the help of https://dcc.ligo.org/LIGO-E1400332. The roundtrip is found to be 24cm which gives an FSR of 1.25GHz.

I use Evan Hall's Python script for calculating the mode spectrum as a function of the cavity length of the metal PMC and overlay the RF sidebands (Green dashed lines) on it (Attachment 18) to check for any HOM coincidence. The width of the lines is the mode splitting due to the cavity astigmatism.

It seems like the only issue might come from a 10th order modes (green ribbon) which are hopefully small enough in reality.

Attachment 1: PMC1Side.jpg
PMC1Side.jpg
Attachment 2: PMC1Front.jpg
PMC1Front.jpg
Attachment 3: PMC1Back.jpg
PMC1Back.jpg
Attachment 4: PMC2Side.jpg
PMC2Side.jpg
Attachment 5: PMC2Back.jpg
PMC2Back.jpg
Attachment 6: PMC2Front.jpg
PMC2Front.jpg
Attachment 7: 20200207_123118.jpg
20200207_123118.jpg
Attachment 8: 20200207_123055.jpg
20200207_123055.jpg
Attachment 9: 20200207_122448.jpg
20200207_122448.jpg
Attachment 10: 20200207_122400.jpg
20200207_122400.jpg
Attachment 11: 20200207_122040.jpg
20200207_122040.jpg
Attachment 12: 20200207_122227.jpg
20200207_122227.jpg
Attachment 13: 20200207_122149.jpg
20200207_122149.jpg
Attachment 14: 20200207_123328.jpg
20200207_123328.jpg
Attachment 15: 20200207_123405_HDR.jpg
20200207_123405_HDR.jpg
Attachment 16: PMC1Screw.jpg
PMC1Screw.jpg
Attachment 17: PMC2BadMirror.jpg
PMC2BadMirror.jpg
Attachment 18: homVersusLength.pdf
homVersusLength.pdf
  15197   Fri Feb 7 09:45:03 2020 shrutiUpdateGeneralAM at X end

I took a few AM TF measurements at the X end for which I:

  • Misaligned the ITMX (then re-aligned it)
  • Opened the X green shutter during the measurements and closed it at the end
  • Moved the Agilent from the PSL area to the X end, the delay line and mixer still remains near the PSL area (will move it soon)
  • Took a bunch of TFs

I will post the data soon.

  15196   Fri Feb 7 02:41:28 2020 KojiUpdateGeneraloffice area temperature

Not sure what's wrong, but the workstation desk is freezing cold again and the room temp is 18degC (64degF).

  15195   Fri Feb 7 02:24:24 2020 gautamUpdateLSCSome short notes

[koji, gautam] 

Plots + interpretation tomorrow.

  • CM_Slow path can be used to stabilize the arm powers somewhat but the AO crossover remains out of reach.
  • The REFL11 (=CARM_B) path offset has to be manually determined - we found that it can change by ~20% depending on the alignment, which maybe isn't surprising given that the mode shapes seen at POP, REFL and AS look like Rorschach inkblots.
  • We saw TRX/TRY regularly hit ~150, and at times even 200 (= recycling gain of ~10). Though any conclusive statement about the PRG can only be made once the lock is stabilized.
  • I was able to take a few CARM loop TFs with an SR785 hooked up at 1Y2. Despite ramping up the AO gain, we saw no effect at high frequencies in the TF shape (the phase bubble continued to roll off at ~100 Hz and there was no visible phase lead even as the AO gain was increased). It has to be estimated what the expected crossover gain is from the experiment with the high BW POY locking (taking into account the net difference in optical gain between POY for single arm and REFL for the full IFO).
  • The fact that I was able to hold the high BW POY lock makes me think that the IMC servo board's IN2 input (and indeed the rest of the IMC locking loop) is functioning as expected. But maybe this board will benefit from a detailed checkout like Koji did for the CM board.

Getting closer... To facilitate this work, I made some convenience scripts that can be run from the CM MEDM screen.

  15194   Thu Feb 6 21:54:13 2020 JonUpdatePSLc1psl bench testing complete

Today I engineered the last piece of the new c1psl system: the multi-bit binary output (mbbo) channels that control the MC servo board gains. These 6-bit channels have to be split across two 4-bit Acromag registers. To enforce synchronous switching, I adapted the latch.py script developed by Gautam to address this problem in c1iscaux. Analogously to the c1iscaux implementation, I scripted the code to automatically run as a systemd service which is launched by the main modbusIOC service. I tested this all using the DB37 LED test board and confirmed it to work.

This now completes the electronics bench testing.

There are still several DB37 connectors to be wired, which carry only spare channels for future use and are not interfaced with the EPICS IOC. Jordan and I discussed this today and he or Chub will complete it shortly. To allow time for the spare channel wiring to be completed (as well as for more locking progress before interruption), Gautam and I think Monday/Tuesday next week would be the earliest possible window to install the new system.

  15193   Thu Feb 6 16:14:44 2020 ranaUpdateGeneraloffice area temperature

I changed the office area thermostate near Steve's desk from 68F to 73F today. Please do not change it.

If anyone from facilities comes to adjust something, please put the details in the elog on the same day so that we can know to undo that change rather than chase down other drifts in the system.

  15192   Thu Feb 6 01:25:50 2020 gautamUpdateLSCLocking updates

Summary:

I managed to partially stabilize the arm citculating powers - they stay in a region in which the REFL 11 signal is hopefully approximately linear and so I can now measure some loop TFs and tweak the transition appropriately. 

Procedure:

The main change I made tonight was to look at the REFL11 signal as I swept the ALS CARM offset through 0. I found that the maximum arm powers coincided with a non-zero REFL11 signal value (i.e. a small CARM offset was required at the input to the CARM_B filter bank). Not so long ago, I had measured the PM/AM ratio for 11 MHz to be ~10^5 - so it's not entirely clear to me where this offset is coming from. Then, I was able to turn on the integrator (z:p = 20:0) in the CARM_B filter bank while maintaining high POP_DC. At this point, I ramped up the IN2 gain on the IMC servo board (= AO path), and was able to further stabilize the power. 

Attachment #shows this sequence from earlier in the evening. Note that in this state, both ALS and IR control of CARM is in effect. The circulating power is fluctuating wildly - partly this is probably the noisy ALS control path, but there is also the issue of the (lack of) angular control - although looking at the transmon QPDs and the POP QPD signals, they seem pretty stable.

The next step will be to try and turn off the ALS control path. Eventually, I hope to transition DARM control to AS55 as well. But at this point, I can at least begin to make sense of some of the time series signals, and get some insight into how to improve the lock.

Quote:

No systematic diagnosis scheme comes to mind.

Attachment 1: semiStableArms.png
semiStableArms.png
Attachment 2: armAngStability.png
armAngStability.png
  15191   Thu Feb 6 01:16:58 2020 gautamUpdateLSCDiagnosis results

Summary:

I did some more detailed tests to see if I could isolate where the excess ALS noise at low CARM offset is coming from, by measuring the spectrum of the IMC error point (in loop). The results, shown in Attachment #1 and #2, are inconclusive.

Details:

Since MC_F didn't show any signatures of elevated noise, I decided to hook up an SR785 to the A excitation bank TEST1 input of the IMC servo board to monitor the in-loop error signal. I initially took a few measurements spanning 800 Hz in frequency, and to my surprise, I found that there was elevated noise in the frequency band we see an increase in the ALS noise, even when the CARM feedback goes to the ETMs (so the IMC cavity is in principle isolated from the main interferometer). This is Attachment #1. So I re-took a couple of measurements (this time only for the case of CARM feedback to the ETMs), with a 200 Hz frequency span, and found no significant noise elevation. This is Attachment #2. I am led to conclude that the IMC error point level changes over time for reasons other than the CARM offset - it'd be nice to have a spectrogram of the IMC error point and compare excursions relative to the median level over a few 10s of minutes, but we don't have this data stream digitized by the CDS system - maybe I will hijack the MC_L channel temporarily to record this data stream. It seems a waste that we're not able to take full advantage of the measured <10pm RMS noise of the IR ALS system.

Attachment 1: IMCspec_ALS.pdf
IMCspec_ALS.pdf
Attachment 2: IMCspec_ALS_smallSpan.pdf
IMCspec_ALS_smallSpan.pdf
  15190   Wed Feb 5 21:13:17 2020 YehonathanUpdateIOOIMC Ringdowns extended data analysis

I translate the results obtained in the previous elog to the IMC 3 mirror cavity. I assume the loss in each mirror in the IMC is equal and that M2 has a negligible transmission.

I find that to a very good approximation the loss per IMC mirror is 2/3 the loss per mirror in the 2 mirror cavity model. That is the loss per mirror in the IMC is 56 ppm. The transmission per mirror in the IMC is the same as in the 2 mirror model, which is 1980 ppm.

The total transmission is the same as in the 2 mirror model and is given by:

\frac{P_0}{P_0+P1}KT^2\approx 90\%

where \frac{P_0}{P_0+P1} is the coupling efficiency to the TEM00 mode.

  15189   Wed Feb 5 21:04:10 2020 YehonathanUpdatePSLBench testing of PSL ai channels

{Yehonathan, Jon}

We retested the failed ai channels. Most of them got fixed by applying the inverse calibration in the yaml file.

We still find some anomalous channels, mostly in the DB25 connector. Turns out, their limits were ill-defined in the EPICS database. Specifying the right limit fixed the issue.

We find one miswired channel (BNC4). We connected the BNC to the right channel on the Acromag unit which fixed the issue.

Overall all the ai channels were successfully bench-tested.

Quote:

I checked the failed channels against the EPICS database definitions and the yaml file inputted to IFOTest. The channels where the readings are something other than +10/0 V, but the high/low values do change, I think can be attributed to one of two things:

  • An incorrect gain and/or offset conversion parameter in the yaml file
  • The EPICS SMOO parameter (smoothing) is set to some long value

I fixed the channel gains/offsets in the master yaml file (PSL.yaml). I also disabled smoothing in the EPICS defintions of the new PSL channels for the purpose of testing. We can uncomment those lines after installing the new chassis if noise is a problem. Please go ahead and re-test the channels again.

Quote:

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

  15188   Wed Feb 5 16:35:12 2020 gautamUpdateLSCDiagnosis plan

The goal is to try and identify the source of the excess ALS noise as the CARM offset is reduced. The idea is to look at the MC_F spectrum (or the IMC error point) in a few conditions:

  1. Regular CARM --> MC2 actuation scheme, PRMI locked on 3f signals, CARM held off resonance.
  2. Regular CARM --> MC2 actuation scheme, PRMI locked on 3f signals, CARM held on resonance.
  3. Alternate CARM --> 7.5*ETMX + 1.5*ETMY, PRMI locked on 3f signals, CARM held on resonance.
  4. Control arms in X/Y basis, lock PRMI on 3f signals and bring the arms into resonance individually, look for excess ALS noise.

#1 vs #2 is like a control experiment, we expect to see the excess noise imprinted on the MC length and hence in MC_F (provided the sensing noise is low enough). #2 vs #3 will be informative of something like backscatter to the PSL increasing the frequency noise. #2/3 vs #4 will help isolate the problem to an individual arm's AUX PDH loop or some optomechanical effect.

I was looking back at some spectra from the last couple of nights but I don't really have an apple-to-apple comparison in the various actuation schemes (some ALS loops were engaged/disengaged), so I'll do a more systematic test tonight. Already, it looks like MC_F is not a good candidate to look for the excess frequency noise, I don't really see a big difference between conditions #1 and #2. According to this, we are looking for an increase at the level of a few 100Hz/rtHz @ ~40 Hz, wheras MC_F is much noisier.

Attachment 1: ALSnoise.pdf
ALSnoise.pdf
  15187   Wed Feb 5 08:57:11 2020 YehonathanUpdatePSLBench testing of PSL ai channels

I checked the failed channels against the EPICS database definitions and the yaml file inputted to IFOTest. The channels where the readings are something other than +10/0 V, but the high/low values do change, I think can be attributed to one of two things:

  • An incorrect gain and/or offset conversion parameter in the yaml file
  • The EPICS SMOO parameter (smoothing) is set to some long value

I fixed the channel gains/offsets in the master yaml file (PSL.yaml). I also disabled smoothing in the EPICS defintions of the new PSL channels for the purpose of testing. We can uncomment those lines after installing the new chassis if noise is a problem. Please go ahead and re-test the channels again.

Quote:

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

  15186   Tue Feb 4 18:13:01 2020 YehonathanUpdatePSLBench testing of PSL ai channels

{Yehonathan, Jon, Jordan}

I tested the ai channels of the new PSL Acromag by looping an already-tested ao channel (C2:PSL-FSS-INOFFSET) back to the different ai channels.

I use Jon's IFOTest with /users/jon/ifotest/PSL.yaml.

I created a spreadsheet for the testing based on the current wiring spreadsheet. I added two columns for the high and low readings for each ai channel (attached pdf).

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

I redid the test on the channel that seemed disconnected to confirm.

I created a yaml file with all the failed channels for retesting called /users/jon/ifotest/PSL_failed_channels.yaml.

Attachment 1: c1psl_wire_testing_-_By_Connector.pdf
c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf
  15185   Tue Feb 4 02:13:02 2020 gautamUpdateLSCLocking updates

Summary:

The CARM-->RF transition remains out of reach. No systematic diagnosis scheme comes to mind.

Details:

  • Config is PRFPMI, SRM is misaligned macroscopically.
  • PRMI can easily be locked with 3f signals while CARM is offset from resonance. Aided by DAFI, I turned on the PR violin filter in the BS output section to prevent it from ringing up, making the lock much more robust.
  • When the CARM offset is reduced
    • POP22 level dips and sometimes goes negative - i don't see this in my simple simulations. POP22 is the trigger signal for MICH/PRCL loops, so to prevent the PRMI lockloss, I mix in some POPDC into the trigger matrix element.
    • Once the circulating power exceeds ~10, the ALS noise apparently increases.
    • The arms "buzz" through resonance, but the power fluctuation is nearly 0-200 in TRX/TRY, corresponding to several CARM linewidths, but all the out-of-loop ALS noise measurements have me believe that we are close to the CARM linewidth in noise. So we should only see ~factor of 2 fluctuation in power.
    • The RF error signal for CARM (=REFL 11) doesn't show any features that i can use to aid the transition / diagnose what is going on systematically.
  • Koji suggested changing the actuation for CARM from MC2 to the ETMs, and check if the MC OSEMs witness the excess motion at small CARM offsets
    • The ALS transition is scripted, so I had to make a modified version that accommodates this changed actuation scheme.
    • The usual CARM-->MC2 matrix element is -1. 
    • The frequency actuation strength of MC2 is ~3x that of the ETMs. Additionally, ETMX has 5x the series resistance of ETMY. So I used the output matrix elements shown in Attachment #1 so as to get the same loop UGF with the same loop gains elsewhere in the chain. Confirmed the actuation strength is the same using the sensing matrix infrastructure and comparing line heights.
    • Attachment #2 shows the measured UGF - both CARM and DARM look okay to me.
  • With this new ALS output matrix actuation scheme, I was able to make it to PRMI + arms on zero offset a couple of times tonight, but the drifting input alignment makes the PRMI lock not so robust anymore.

TBC. Mercifully at least the shaker stayed still tonight.

Attachment 1: modifiedOutMat.png
modifiedOutMat.png
Attachment 2: OLTFs.pdf
OLTFs.pdf
  15184   Mon Feb 3 15:22:39 2020 JonUpdatePSLc1psl progress/Acromag ADC grounding

I tested the c1psl AO channels on the electronics bench on Friday. While I found all the wiring to be correct, some of the channels exhibited excess noise with all appearances of a grounding problem.

Today Jordan, Gautam, and I investigated this further. It is indeed a grounding problem, but actually with the Acromag ADCs. The Acromag DAC outputs are single-ended (return is grounded), so (for the purpose of a loopback test) I would expect to leave the ADC inputs ungrounded. This is the configuration I tested Friday. Today we also tested driving the ADC with a floating source. The ADC noise behavior is exactly the same, whether the source end is grounded or not.

However, grounding the minus pin of the ADC channel eliminates the noise. We don't understand why this seems to be required irrespective of the driving source, so there something we're missing about the ADC design. As it turns out, this same fix was made to the AI channels of the previously-upgraded Acromag machines. I know Chub and I had to do this for the AI channels of c1vac, but at the time we thought the source grounding was causing the issue. However, today Jordan and I looked inside c1iscaux, which Chub wired, and confirmed that its AI channels are wired in the same way.

So in any case, Jordan is grounding the c1psl AI channels in the same way as c1iscaux. Once this is done, we'll continue with the bench testing tomorrow.

gautam: here are my notes about this issue when i was doing the c1iscaux testing. As I note there, "previously-upgraded Acromag machines" in the plural may be a bit of a stretch - I have no idea what the grounding situation is in c1susaux / c1auxex for example.

  15183   Mon Feb 3 13:54:10 2020 YehonathanUpdateIOOIMC Ringdowns extended data analysis

I extended the ringdown data analysis to the reflected beam following Isogai et al.

The idea is that measuring the cavity's reflected light one can use known relationships to extract the transmission of the cavity mirrors and not only the finesse.

The finesse calculated from the transmission ringdown shown in the previous elog is 1520 according to the Zucker model, 1680 according to the first exponential and 1728 according to the second exponential.

Attachment 1 shows the measured reflected light during an IMC ringdown in and out of resonance and the values that are read off it to compute the transmission.

The equations for m1 and m3 are the same as in Isogai's paper because they describe a steady-state that doesn't care about the extinction ratio of the light.

The equation for m2, however, is modified due to the finite extinction present in our zeroth-order ringdown.

Modelling the IMC as a critically coupled 2 mirror cavity one can verify that:

m_2=P_0KR\left[T-\alpha\left(1-R\right)\right]^2+\alpha^2 P_1

Where P_0 is the coupled light power 

P_1 is the power rejected from the cavity (higher-order modes, sidebands)

K=\left(\mathcal{F} /\pi \right )^2 is the cavity gain.

R and T are the power reflectivity and transmissivity per mirror, respectively.

\alpha^2 is the power attenuation factor. For perfect extinction, this is 0.

Solving the equations (m1 and m3 + modified m2), using Zucker model's finesse, gives the following information:

Loss per mirror = 84.99 ppm
Transmission per mirror = 1980.77 ppm
Coupling efficiency (to TEM00) = 97.94%
Attachment 1: IMCTransReflAnalysis_anotated.pdf
IMCTransReflAnalysis_anotated.pdf
  15182   Fri Jan 31 16:57:09 2020 gautamUpdateGeneralMetal PMC parts

Jon brought over a box of parts for constructing the metal PMCs. I have stored it along the Y-arm, on top of the green optics cabinet.

I didn't do an exhaustive inventory check, but the following are the rough contents of the box:

  • 41 deg AoI flat mirrors, R=99% @ 1064nm --- 11 pcs
  • 6.8 deg AoI curved mirrors --- 5 pcs
  • PZTs --- 3pcs
  • Metal PMC body --- 2 pcs
  • "Baked PZT endcaps" --- 3 pcs
  • Ball bearings, clamps, misc hardware

I didn't inspect the optics but since we have so many, I am hoping we can find 3 good quality ones for one cavity at least. We should check that the geometry is suitable for our RF sideband frequencies.  

  15181   Fri Jan 31 16:04:30 2020 gautamUpdateIOOInput pointing drift

One factor which hampers locking efforts is the apparent drift of the input beam into the IFO. Over timescales of ~1 hour, I have noticed that the spot on the AS camera drifts significantly (~1 spot size) in pitch. The IPPOS QPD bears out this observation, see Attachment #1. The IMC WFS control signals do not show a correlated drift, hence my claim that the TTs are to blame.

I am able to correct this misalignment by moving TT1 in pitch (see Attachment #2, which shows some signals from a ~1 hour PRMI lock, during which time the pointing drifted, and I corrected it by moving TT1 pitch). Assuming the problem is purely TT1 pitch drifting, this corresponds to 3mm / 6m ~500urad of shift in 1 hour - seems very large. The fact that the drift is only present in pitch and doesn't really show up in yaw makes me think the problem is likely mechanical (unless the voltage to the top two coils is drifting relative to the bottom, but no LR drift, which would be very coincidental). At the moment, this is just an annoyance, but it'd be good for this problem to be fixed.

I haven't yet figured out how to make ndscope export these plots to SVG preserving the dark color theme, hence the weird light axes...

Attachment 1: IPdrift.pdf
IPdrift.pdf
Attachment 2: IPdrift_PRMI.pdf
IPdrift_PRMI.pdf
  15180   Thu Jan 30 22:02:42 2020 shrutiUpdateGeneraldelay line frequency discriminator for PM

I could not find any level 3 mixers, but by adjusting the beat frequency the power in each of the delay line channels rose to almost 6.5 dBm.

In short: Delay line seems to work

Things I did earlier today:

  1. Played with the slow servo on the FSS screen, but then reset the parameters to what was there before (Later found out that this was to lock the PSL freq to the IMC when the IMC power is significant.)
  2. Connected the AG 4395A to the X PZT
  3. Closed the PSL shutter

Transfer function measurement: (Refer Attachment 1)

Everything about the setup remained as I had left it earlier: described in elog 15174

except

  • SR560 gain set to 10, DC coupled
  • DC block at channel A of Agilent (The measurement was A/R)

I did not use a slow servo, but took individual sweeps adjusting the PSL temperature each time to bring the error voltage between +/-25 mV. The beat frequency was over 100 MHz.

For the plot posted in Attachment 1, the measurement paramters are the following. Will do further measurements/analysis tomorrow.

# AG4395A Measurement - Timestamp: Jan 30 2020 - 21:58:00
# Parameter File: TFAG4395Atemplate.yml
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 50000.0, 50000.0
# Stop Frequency (Hz): 1000000.0, 1000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 1
# Auto Bandwidth: Off, Off
# IF Bandwidth: 1000.0, 1000.0
# Input Attenuators (R,A,B): 0dB 0dB 0dB 
# Excitation amplitude = -20.0dBm

Quote:

yes, its fine to use this with a level 3 or level 7 mixer; let's see some PM transfer functions !

Quote:

Is this sufficient enough for the mixer to work?

Attachment 1: Figure_2.png
Figure_2.png
  15179   Thu Jan 30 17:41:10 2020 gautamUpdatePSLErrant FSS_INOFFSET change

You can trend the data for the past few hours and see what the appropriate value. I think these tests should only be done when whoever is running a test is in the lab.

P.S. I was surprised that the IMC didn't lose lock when this step was applied. I manually stepped this voltage between +/- 10 V and didn't see any response in the FSS readbacks. Either the channel doesn't work, or there is a divide by 40 in the physical circuit or something...

Quote:

A script I was testing errantly set C1:PSL-FSS_INOFFSET => 10 V at about 5:30 pm. I manually reverted the channel value to 0, but I don't know what the value was initially. Someone please check this value if there are problems locking the FSS.

  15178   Thu Jan 30 17:31:28 2020 JonUpdatePSLErrant FSS_INOFFSET change

A script I was testing errantly set C1:PSL-FSS_INOFFSET => 10 V at about 5:30 pm. I manually reverted the channel value to 0, but I don't know what the value was initially. Someone please check this value if there are problems locking the FSS.

  15177   Thu Jan 30 15:24:10 2020 ?UpdateGeneraldelay line frequency discriminator for PM

yes, its fine to use this with a level 3 or level 7 mixer; let's see some PM transfer functions !

Quote:

Is this sufficient enough for the mixer to work?

  15176   Thu Jan 30 12:52:10 2020 JonUpdateBHDMetal OMCs procured

Last night Yehonathan and I located the two steel PMCs in the QIL, with help from Anchal. They are currently sitting on my desk in Bridge, inside a box that also contains optics and other OMC parts. I will bring them over to the 40m the next time I come.

  15175   Wed Jan 29 12:40:24 2020 YehonathanUpdateIOOIMC Ringdowns preliminary data analysis

I analyze the IMC ringdown data from last night. 

Attachment 1 shows the normalized raw data. Oscillations come in much later than in Gautam's measurement. Probably because the IMC stays locked.

Attachment 2 shows fits of the transmitted PD to unconstrained double exponential and the Zucker model.

Zucker model gives time constant of 21.6us

Unconstrained exponentials give time constants of 23.99us and 46.7us which is nice because it converges close to the Zucker model.

Attachment 1: IMCRingdownNormalizedRawdata.pdf
IMCRingdownNormalizedRawdata.pdf
Attachment 2: IMCTransPDFits.pdf
IMCTransPDFits.pdf
  15174   Wed Jan 29 12:29:33 2020 shrutiUpdateGeneraldelay line frequency discriminator for PM

 Today I began working on a TF measurement based on the delay line frequency discriminator setup in elog 4254 using a single mixer (without the 'I' and 'Q' readout).

For this, I re-organised the setup for the PLL measurement of the transfer function (elog 15148), increasing the HEPA for the initial changes while the PSL door was open, and then reverting it back to ~30%:

  • I removed the 20dB coupler and connected the splitter directly after the amplifer to split the beat note signal into two coaxial cables one of which was ~1.5m longer than the other
  • The recombined signals were combined in a mixer outside the PSL enclosure. I also replaced the 1.9 MHz LPF with a 5 MHz LPF.
  • I used an SR 560 to amplify the signal after the LPF.

With the above setup the power that was seen at each channel of the delay line was <1dBm, which is not ideal for the any of the available mixers.

After the group meeting, I changed the amplifer to ZHL-3A (that is near the beat mouth) instead of a ZFL-500HLN because it had a higher gain (~28dB as opposed to ~19dB of the latter). The power seen at each of the delay line channels is over 5.5 dBm. This is consistent with the estimation 0 dBm beat -> -20 dBm after 20dB coupler -> 8 dBm after amplifier -> 5 dBm after splitter with insertion loss of 3 dB.

Is this sufficient enough for the mixer to work? In Attachment 1: A shows the mixer output (point B in Attachment 2) when the IMC is locked, in B the IMC is unlocked at the middle of the spectrum, and each of the dips show the DC voltage being sent to the PSL temperature servo being decreased by 0.01 V.

Gautam pointed me to the location of a few other RF amplifiers (ZHL-32A+, ZHL-1A) which don't possess a higher gain but can be used without disrupting the ALS related work (I was told).

For shorter duration changes that I made later, I opened and closed the PSL enclosure doors without changing the HEPA.

Attachment 2 shows the current setup as is, but I might add a PSL servo tomorrow to stabilise its frequency corresponding to a null mixer output without driving anything else.

Attachment 1: 20200128.png
20200128.png
Attachment 2: IMG_BB01C068495A-1.jpeg
IMG_BB01C068495A-1.jpeg
  15173   Wed Jan 29 03:05:47 2020 rana, gautamUpdateSUSMC misalignments / sat box games

In the last couple days, as the IMC ringdowns have been going on, we have noticed that the MC is behaving bad. Misaligning, drifting, etc.

Gautam told me a horror story about him, Koji, and melted wires inside the sat boxes.

I said, "Its getting too hot in there. So let's take the lids off!"

So then we:

  1. Removed the lid (only 4 screws were still there)
  2. cut off some of the shield - ground wires and insulated them with electrical tape
  3. squished the IDC connectors on tightly
  4. left it this way to see if MC would get better - certainly the painfully hot heatinks inside the box were now just 110 F or so

After some minutes, we saw no drifting. So maybe my theory of "hot heatsink partially shorting a coil current to GND through partially melted ribbon cable" makes sense? IF this seems better after a month, lets de-lid all the optics.

Let's look at some longer trends and be very careful next to MC2 for the next 3 days! I have put a dangerous mousetrap there to catch anyone who walks near the vacuum chamber.

gautam: the grounding situation per my assessment is that the shield of all the IDC cables are connected to a common metal strip at 1X5 - but in my survey, I didn't see any grounding of this strip to a common ground.

Attachment 1: IMG_8366.JPG
IMG_8366.JPG
  15172   Wed Jan 29 00:29:43 2020 gautamUpdateLSClocking 2020

The goal tonight was to go through the locking scripts to see if I could recover the state from November 2019, when I could have the arm lengths controlled by ALS, and sit at zero CARM offset with the PRMI remaining locked and the arm powers fluctuating between 0-300. The IR-->ALS transitions went smoothly tonight, and the PRMI locking was also fairly robust when the CARM offset was large, but was less good when reduced to 0. Nevertheless, it is good to know that the system can be restored to the state it was late last year. Next step is to figure out how to keep the PRMI locked and get the AO path engaged, this was what I was struggling with before the new year.

Attachment 1: PRFPMI_2020Jan.png
PRFPMI_2020Jan.png
  15171   Wed Jan 29 00:27:13 2020 gautamUpdateComputer Scripts / Programsmcup / mcdown modified

To fix the apparent slowness of execution of the caput commands on megatron, I changed the "ewrite" macro in the mcup and mcdown scripts to use ezcawrite instead of caput. The old lines are simply commented out, and can be reverted to at any point if we so desire. After these changes, we saw that both scripts complete execution much faster.

  15170   Tue Jan 28 20:51:37 2020 YehonathanUpdateIOOIMC WFS servos stable again

I resume my IMC ringdown activities now that the IMC is aligned again.

To avoid any accidental misalignments Gautam turned off all the inputs to the WFS servo.

I set up a PD and a lens as in attachment 1 (following Gautam's setup).

I connect the REFL, TRANS and INPut PDs to the oscilloscope.

I connect a Siglent function generator to the AOM driver. I try to shut off the light to the IMC using 1V DC waveform and pressing the output button manually. However, it produced heavily distorted step function in the PMC trans PD.

I use a square wave with a frequency of 20mHz instead with an amplitude of 0.5V offset of 0.25V and dutycycle of 1% so there will be minimal wasted time in the off state. I get nice ringdowns (attachment 2) - forgot to take pictures. The autolocker slightly misaligns the M2 every time it is acting, so I manually align it everytime the IMC gets unlocked.

Data analysis will come later.

I remove the PD and lens and reenable the WFS servo inputs. The IMC locks easily. The WFS outputs are very different than 0 now though.

  15169   Tue Jan 28 19:40:15 2020 shrutiUpdateGeneralPLL / PM measurement of Xend NPRO PZT

Over the past few days, I have been trying to make measurements of the phase modulation transfer function by modulating the X end laser PZT via PLL.

The setup was modified every time during the experiment in the same manner as mentioned in elog 15148.

I could not make the PLL lock for long enough to take a proper TF measurement, resulting in TFs that look like Attachment 1. The next step would be to use the method of a delay line frequency discriminator instead of the PLL.


Comments about locking with LB1005 PI controller:

  1. I do not understand why the high PI corner frequency of 1kHz or 3kHz was required to lock.
  2. The rms level of the error signal when locked was ~100 mV, which is 25% of the total mixer range (~400 mVpp). Decreasing the gain only caused the loop to go out of lock and did not decrease this noise in the error signal.
  3. The setup was also partly inside the PSL enclosure, with the HEPA turned to 100%, which is probably a noisy environment for this measurement. Closing and opening the shutters or any disturbance near the enclosure resulted in movement of the beat note up to 5 MHz.
  4. It may have been a better idea to actuate the PSL laser instead of the X NPRO because of its larger range, but would this solve the issue with the noise?
Attachment 1: PMTF.pdf
PMTF.pdf
Attachment 2: BeatSpectrum.pdf
BeatSpectrum.pdf
  15168   Tue Jan 28 19:12:30 2020 JonConfigurationPSLSpare channels added to c1psl chassis

After some discussion with Gautam, I decided to build more spare channels into the new c1psl machine. This is anticipation of adding new laser and ISS channels in the near future, to avoid having to disconnect the installed chassis and pull it out of the rack. The spare channels will be wired to DB37M feedthroughs on the front side of the chassis, with enough wire length to be able to pull the breakout boards out of the front to reconfigure their wiring as needed (e.g., split off channels onto a separate connector).

To have enough overhead, this will require installing 1 additional ADC unit (XT1221) and 1 additional DAC (XT1541). We have enough spare BIO channels among the existing units (both sinking and sourcing). This will give us:

  • 13 spare ADC channels
  • 14 spare DAC channels
  • 16 spare sinking BIO channels
  • 12 spare sourcing BIO channels

The updated c1psl chassis wiring assignments are attached. It adds 4 new DB37M connectors for the spare channels (highlighted in yellow) and fixes one typo Jordan found while wiring today. The most current spreadsheet is available here.

Attachment 1: c1psl_feedthrough_wiring_v2.pdf
c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf
  15167   Tue Jan 28 17:36:45 2020 gautamConfigurationComputersLocal EPICS7.0 installed on megatron

[Jon, gautam]

We found that the caput commands were taking much longer to execute on megatron than on pianosa (for example). Suspecting that this had something to do with the fact that megatron was using EPICS binaries from the shared NFS drive which were compiled for a much older OS, I installed the latest stable release of EPICS on megatron. The new caput commands execute much faster. I also added the local EPICS directory to the head of the $PATH variable used by the MC autolocker and FSS Slow scripts, so that they use the new caput command. But mcup is still slow - maybe my new path definition isn't picked up and it is still using the NFS binaries? To be looked into...

Quote:

There were a bunch of medm processes stalled on megatron (connected with screenshot taking). To see if they were interfering with the other scripts, I killed all of the medm processes, and commented out the line in the crontab that runs the screenshots every 10 mins. Let's see if this improves stability.

  15165   Tue Jan 28 16:01:17 2020 gautamUpdateIOOIMC WFS servos stable again

With all of the shaking (man-made and divine), it was a hard to debug this problem. Summary of fixes:

  1. The beam was misaligned on the WFS 1 and 2 heads, as well as the MC2 trans QPD. I re-aligned the former with the IMC unlocked, the latter (see Attachment) with the IMC locked (but the MC2 spot centering loops disabled).
  2. I reset the WFS DC and RF offsets, as well as the QPD offsets (once I had hand-aligned the IMC mirrors to obtain good transmission).

At least the DC indicators are telling me that the IMC locking is back to a somewhat stable state. I have not yet checked the frequency noise / RIN.

Attachment 1: QPD_recenter.png
QPD_recenter.png
  15164   Tue Jan 28 15:39:04 2020 gautamConfigurationComputersSluggish megatron?

There were a bunch of medm processes stalled on megatron (connected with screenshot taking). To see if they were interfering with the other scripts, I killed all of the medm processes, and commented out the line in the crontab that runs the screenshots every 10 mins. Let's see if this improves stability.

  15163   Tue Jan 28 14:33:24 2020 gautamUpdatePSLInferred free-running frequency noise

To conclude my PMC noise investigations: Attachment #1 shows the PMC noise inferred from the calibrations earlier in this thread and the fitted OLTF for the PMC loop. Attachment #2 compares the frequency noise (inferred from the error point of the PMC servo) when the IMC is locked / unlocked. I don't know what to make of the fact that the PMC suggests improvement from ~20 Hz onwards already - does this mean that the NPRO noise model is wrong by 1 order of magnitude at 30 Hz?

  • The IMC was locked for the measurement shown in Attachment #1.
  • The in-loop spectra of the error (at the I/F output of the mixer) and control (at TP3) signals were measured with the SR785.
  • The control signal voltage monitors don't seem to work - neither the front panel LEMO nor the signals hooked up to the CDS system show me sensible shapes for the spectra between 1-3 Hz.
  • To convert in loop to free-running, I multiplied the measured error (control) signal spectra by \left | 1-L \right | (\left | \frac{L}{1-L} \right |), where L is the OLTF. THe control signal was pre-processed by multiplying by a pole at 11.3 Hz, corresponding to the LPF formed by the 63.3 kohm series resistor and the 225 nF PZT capacitance.
  • The "NPRO noise model" curve is 10^4/f Hz/rtHz.

While I initially thought the 1/f^2 rise below ~100 Hz is attributable to the IMC cavity length fluctuations, I found that this profile is present even in the measurement with the PSL shutter closed. I am not embarking on a detailed PMC noise budgeting project for now. Note however that we are not shot noise limited anywhere in this measurement band.

  • The measured TF (dots in Attachment #5) was fit with LISO (solid lines in Attachment #5) to allow inferring the out-of-loop servo noise by monitoring the in-loop noise (that plot to follow).
Attachment 1: inLoopNoise_IMClocked.pdf
inLoopNoise_IMClocked.pdf
Attachment 2: freqNoiseComparison.pdf
freqNoiseComparison.pdf
  15162   Tue Jan 28 08:26:53 2020 ranaFrogsPEMshaking

https://breakthrough.caltech.edu/magazine/2019-aug/#article-Listening-with-Light

ELOG V3.1.3-